OpenStudy (anonymous):

use the magnitude (sqrt 113) and the angle (131.19 degrees) of the vector v to write it in the form (|v|cos theta,|v|sin theta).

OpenStudy (anonymous):

oh, the original vector is: v=(-7,8), sorry!

OpenStudy (johnweldon1993):

Well you already have all the ingredients <|v|cos theta , |v|sin theta> So we just have \[\large <\sqrt{113}\cos 131.19 , \sqrt{113}\sin 131.19>\]

OpenStudy (anonymous):

tbank you! would that be the final answer?!